OpenStudy (anonymous):

A source emits sound waves isotropically. The intensity of the waves 2.50 m from the source is 1.91 x 10^-4 W/m^2. Assuming the energy waves are conserved, the power of the source is?

OpenStudy (anonymous):

well I havent done in this in more than a year. but from the units it looks like the formula for intensity = Power / Area ( and I dont think there is a proportionality constant lying around ).

OpenStudy (anonymous):

pellet

OpenStudy (anonymous):

I put volume instead of area lol

OpenStudy (anonymous):

The area of a sphere is \[4 \pi r^{2}\]so the power is \[IA=I \times4 \pi r^{2}\]

OpenStudy (anonymous):

thats why someone is replying

OpenStudy (anonymous):

\[P = 4(1.91 \times 10^{-4} ) (2.5^2)\pi \]

OpenStudy (anonymous):

i used area of circle instead of sphere, thanks for correcting me!!